OpenStudy (anonymous):

the cost in cents to produce x cups of Mountain Thunder Lemonade at Junior's Lemonade stand is C(x)=18x+240 x>=0 and the price function in cents per cup is p(x) =90-3x 0<=x<=30 find the price to charge per cup in order to maximize profit and i found 12 cups need to be sold to maximize profit

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Does the attachment show the same problem as yours? @ninab731

OpenStudy (anonymous):

p(x) =\[-3x^2+72x-240\] Maximum profit = \[\frac{ -72 }{ -6 }=12\] \[-3(12)^2+72(12)-240=192 \] $1.92 maximum profit yes that attachment is correct

Directrix (directrix):

Here are the instructions.

OpenStudy (anonymous):

@Directrix http://awesomescreenshot.com/0df1ps169d

Directrix (directrix):

I will have to read an example of this type problem and how to solve it. $1.92 for maximum profit seems low.

OpenStudy (anonymous):

http://awesomescreenshot.com/0c11ps1fd7

Directrix (directrix):

I don't follow how the reasoning behind getting 12 cups for maximum profit. Will you explain that in words?

OpenStudy (anonymous):

using the quick vertex formula -b/2a so -72/-6 gives us 12

OpenStudy (anonymous):

then plugged that into the equation to get $1.92

Directrix (directrix):

To break even, you have to sell 4 cups or 20 cups. That comes from finding the zeros of the function f(x) = -3x² + 72x - 240 in the interval [0,30].

Directrix (directrix):

Okay, you got the maximum number of cups from the vertex of f(x). Is that correct?

OpenStudy (anonymous):

yes thats right

Directrix (directrix):

Okay, I get the 12 cups. So, that's 192 cents which is $1.92.

OpenStudy (anonymous):

right so how do we find the price to charge per item in order to maximize profit

Directrix (directrix):

Where is the cost function?

OpenStudy (anonymous):

\[C(x)=18x+240 \]

Directrix (directrix):

C(0 cups)=18*0+240 where x>=0 C(0) = $2.40 so it costs $2.40 before you can make any cups of lemonade.

OpenStudy (anonymous):

ok so where are they getting it as saying the price per cup should be set at 54 cents to maximize profit

Directrix (directrix):

What is the cost for 12 cups?

OpenStudy (anonymous):

oh ok i get it so plug the 12 for C(12)?

Directrix (directrix):

I'm just asking as I think. I don't know for sure doing that will give 54 cents. Let me think.

OpenStudy (anonymous):

18(12)+240 = 456

Directrix (directrix):

Look at this: the price function in cents per cup is p(x) =90-3x 0<=x<=30 p(12 cups) = 90 - 3*12 = 90 - 36 = ? @ninab731

OpenStudy (anonymous):

ah 54!!!

Directrix (directrix):

This is a crazy problem and no way would I do all that work to make $1.92 at the lemonade stand.

OpenStudy (anonymous):

haha thanks a bunch for the help

Directrix (directrix):

You did all the work. It was fun.

OpenStudy (anonymous):

be a lot of work to make a 1.92 haha no way lol